OpenStudy (anonymous):

What is the angular velocity of a point on the tread of a 15-inch diameter tire on a car travelling 55 mph? Answer in radians/sec.

OpenStudy (anonymous):

This is a simple unit conversion problem. In order to solve you have to know that the speed of rotation directly corresponds to the the speed of the car, since the tire is not slipping. \[\frac{55mi}{1hr}*\frac{1hr}{3600s}*\frac{5280ft}{1mi}*\frac{12inches}{1ft}*\frac{1rad}{7.5inches}=x\frac{rad}{s}\]

OpenStudy (anonymous):

thanks!